OpenStudy (anonymous):

get dy/dx from y = sin^-1 x + cos^-1 x

OpenStudy (anonymous):

0

OpenStudy (anonymous):

\[{dy \over dx} \sin^{-1} x = {1 \over \sqrt{1 - x ^{2}}}\] and \[\[{dy \over dx} \cos^{-1} x = -{1 \over \sqrt{1 - x ^{2}}}\]\]